# The Impact of Baryonic Effects on the Dynamical Masses inferred using Satellite Kinematics

### v1

@@ 1 Introduction
The advent of wide-field galaxy surveys has enabled data-driven approaches to empirically infer the galaxy-halo connection and its evolution over time. Popular methods include galaxy clustering (e.g. Berlind & Weinberg 2002; Yang et al. 2003; van den Bosch et al. 2007; Zehavi et al. 2011; Hearin & Watson 2013; Guo et al. 2015a, b, 2016; Zentner et al. 2019), galaxy–galaxy lensing (e.g. Guzik & Seljak 2002; Mandelbaum et al. 2006, 2016; Sonnenfeld & Leauthaud 2018; Leauthaud et al. 2012, 2017), and subhalo abundance matching (SHAM) (Kravtsov et al. 2004; Conroy et al. 2006; Vale & Ostriker 2006; Reddick et al. 2013). We refer to Wechsler & Tinker (2018) for a comprehensive review of the different approaches to probe and understand the galaxy-halo connection.
@@
A less commonly used technique for probing the galaxy-halo connection is satellite kinematics. This method represents the earliest approach, dating back to the first evidence for dark matter by Zwicky (1933). It involves measuring the line-of-sight velocities of satellite galaxies relative to their central galaxy. Since satellite galaxies are assumed to trace the dark matter potential wells, their dynamics can be used to constrain the gravitational potential and mass of the host halo. Although some concerns have been raised about this method (for an extensive review, see Lange et al. 2019a, and references therein), numerous studies have progressively refined satellite kinematics and demonstrated its utility as a robust tool for constraining the galaxy–halo connection (van den Bosch et al. 2004; More et al. 2009b, 2011; Lange et al. 2019a, b; Mitra et al. 2024; van den Bosch et al. 2019). In particular, Basilisk, a Bayesian hierarchical inference formalism, introduced by van den Bosch, Lange & Zentner (2019) and refined by Mitra et al. (2024), has demonstrated exceptional power to probe the galaxy-halo connection using the kinematics of satellite galaxies in large redshift surveys such as the Sloan Digital Sky Survey (SDSS; York et al. 2000). Basilisk  forward models the 2-D projected phase-space distribution of satellite and central galaxy pairs. A key advantage of this forward-modelling approach is its ability to directly compute the likelihood of the full projected phase-space data given the model without the need for stacking the data in luminosity bins. Consequently, it can simultaneously solve for halo mass, the radial profile of the satellite galaxies and orbital anisotropy of the satellite galaxies, while properly accounting for scatter in the galaxy–halo connection. In addition, it properly accounts for biases and selection effects in the data, such as incompleteness, fibre collisions, and interlopers (Lange et al. 2019a; van den Bosch et al. 2019; Mitra et al. 2024). Recently, Mitra et al. (2024) applied Basilisk  to SDSS-DR7 data, yielding inferences on the galaxy–halo connection that are consistent with, yet tighter than, previous constraints derived from other methods like galaxy clustering and galaxy–galaxy lensing, while also overcoming the mass-anisotropy degeneracy and without being sensitive to halo assembly bias.
@@
An important assumption in satellite kinematics studies is the halo’s density profile, which shapes the gravitational potential. Often, such studies, including those conducted with Basilisk, assume that halos are composed solely of dark matter, neglecting the presence of baryons. However, baryons, in the form of gas and stars, can play a critical role in shaping the structure and dynamics of dark matter halos. Processes such as gas cooling, star formation, and feedback from supernovae or active galactic nuclei (AGN) can modify the distribution of both baryons and dark matter. Early models of baryonic effects suggested that gas cooling and dissipation, followed by star formation, can deepen the gravitational potential well (White & Rees 1978; White & Frenk 1991) and lead to adiabatic contraction of dark matter (Blumenthal et al. 1986). With the advent of cosmological hydrodynamical simulations, detailed studies have further examined the role of baryons in shaping halo structure, clustering, and abundance (e.g. Duffy et al. 2010; Schaller et al. 2015a, b; Sorini et al. 2024; Beltz-Mohrmann & Berlind 2021; Euclid Collaboration et al. 2024). These effects are highly complex, as the influence of baryonic physics depends on factors such as halo mass, environment, and redshift. Moreover, the impact of baryons is extremely sensitive to the specific feedback prescriptions, which differ significantly across hydrodynamical simulations (see review by Somerville & Davé 2015). Despite these challenges, hydrodynamical simulations are often used for correcting for “baryonic effects", that is deviations from dark-matter-only models, when constraining for example the matter power spectrum (e.g. White 2004; van Daalen et al. 2011; Schaller et al. 2024; Schaye et al. 2023; Hernández-Aguayo et al. 2023; Schaller et al. 2024) and weak lensing observables (Broxterman et al. 2024; Semboloni et al. 2011).
@@
An alternative approach is to quantify baryonic effects analytically, providing a flexible tool to complement hydrodynamical simulations and systematically test their impact on cosmological and large-scale structure inferences. While such an approach has been used to study baryonic effects on the matter power spectrum and weak lensing (Fedeli 2014; Fedeli et al. 2014; Semboloni et al. 2011; Mead et al. 2015; Schneider & Teyssier 2015; Schneider et al. 2019), to our knowledge baryonic corrections have not yet been applied to the inference from satellite kinematics. The main goal of this paper is to rectify this situation and to study how baryons impact the line-of-sight velocity dispersion of satellite galaxies modelled as tracers of the underlying gravitational potential well. We do so by comparing the results of a dark-matter-only (DMO) model with one in which the halo hosts a central galaxy and an extended distribution of diffuse stars and gas. Crucially, we examine how the ejection of baryons from the host halo due to feedback processes impacts the potential well, and consequently, the satellite kinematics. We use these results to develop correction functions that can be used to correct the line-of-sight velocity dispersions of satellites computed using a DMO model for the presence of baryons. We also explore conservative bounds on this baryonic correction model that can be used to marginalize over uncertainties arising from incomplete understanding of physical processes associated with galaxy formation. We specifically tailor these results so they can be used in Basilisk  when applied to SDSS data. In a companion paper (Mitra et al. in prep), it is shown that these baryonic corrections, which are typically only at the level of a few percent, can have an appreciable impact on the inference, affecting both the inferred galaxy halo connection as well as cosmological constraints.
@@
Throughout this paper, we adopt the flat Planck18 CDM cosmology (with BAO constraints, Table 2 in Planck Collaboration VI 2020), with $H_{0}$=67.66 km s-1 Mpc-1, $\Omega_{\rm m,0}$ = 0.3111, $\Omega_{\rm b,0}$ = 0.0490, $n_{\rm s}=0.9665$ and $\sigma_{8}=0.810$.
@@ 2.1 Satellite kinematics
$\sigma_{\rm r}^{2}(r)=\frac{G}{r^{2\beta}\,n_{\mathrm{sat}}(r)}\int_{r}^{\infty}r^{\prime 2\beta-2}n_{\mathrm{sat}}(r^{\prime})\,M_{\rm tot}(<r)\,\mathrm{d}r^{\prime}\,$, (1)
(van den Bosch et al. 2019), where $M_{\rm tot}(<r)$ is the total enclosed mass profile, $n_{\mathrm{sat}}(r)$ is the radial number density profile of the satellites, and
@@
Integrating the velocity dispersion in the line-of-sight direction along the line-of-sight, weighted by the number density of the tracer population, yields the line-of-sight velocity dispersion
@@
Throughout we follow Mitra et al. (2024) and assume that the radial number density distribution of satellite galaxies follow a generalized Navarro, Frenk & White (1996) (NFW) profile given by
@@
$c_{200}$ Concentration of the halo derived from concentration-halo mass relation. Diemer & Joyce (2019)
@@
$f_{\mathrm{*,diffuse}}$ Fraction of mass in the diffuse stellar component. Best-fit derived from EAGLE
@@
$f_{\mathrm{eject}}$ Mass fraction ejected from halo. Best-fit derived from EAGLE
@@ 2.2 Mass models
The key ingredient for computing $\sigma_{\rm ap}$ is a model for the total enclosed mass, $M_{\rm tot}(<r)$, which enters the expression for the radial velocity dispersion given by equation (1). Here we describe the two mass models that we use for our investigation: a DMO model and a baryonic model.
@@
$M_{\rm tot}(<r)=$ $M_{\rm DM}(<r)+M_{\rm\ast,cen}(<r)+\,$
$M_{\rm\ast,diffuse}(<r)+M_{\rm gas}(<r)\,.$ (9)
We further define $M_{\rm tot}\equiv M_{\rm tot}(<r_{200})$ as the total bound mass of the halo in the baryonic model, which is related to the DMO mass according to
@@ 2.2.1 Dark Matter
with $c_{200}=r_{200}/r_{\rm s}$ the halo concentration. Throughout we compute this concentration parameter using the redshift-dependent concentration-mass relation of Diemer & Joyce (2019).
@@ 2.2.2 Central Galaxy
$M_{\mathrm{\ast,cen}}(<r)=M_{\mathrm{\ast,cen}}\frac{r^{2}}{(r+a)^{2}}\,.$ (16)
@@ 2.2.3 Diffuse stellar component
Whereas $f_{\mathrm{\ast,cen}}(M_{\rm DMO})$ is tightly constrained via the SHMR, the baryonic mass fraction of the diffuse stellar component, $f_{\mathrm{\ast,diffuse}}$, is poorly constrained. The main reason is that stellar halos typically have extremely low surface brightness, making them difficult to detect. In addition, the outer light profile of the central galaxy typically blends smoothly into that of the stellar halo, hampering a unique decomposition, both in observations (e.g., Huang et al. 2018) and in numerical simulations (e.g., Pillepich et al. 2018). Despite these challenges, it has become clear that $f_{\mathrm{\ast,diffuse}}$ increases with halo mass, albeit with considerable scatter (Merritt et al. 2016; Deason et al. 2019; Ragusa et al. 2023; Montes & Trujillo 2019). We describe this mass dependence using the following simple analytical form:
$f_{\mathrm{\ast,diffuse}}(M_{\rm DMO})=\frac{f_{\mathrm{\ast,diffuse,max}}}{2}\left[1+{\rm erf}\left(\log\frac{M_{\rm DMO}}{M_{\mathrm{crit}}}\right)\right]\,$, (17)
which transitions from $f_{\mathrm{\ast,diffuse}}=0$ for $M_{\rm DMO}\ll M_{\rm crit}$ to a maximum of $f_{\mathrm{\ast,diffuse,max}}$ when $M_{\rm DMO}\gg M_{\rm crit}$. We treat $f_{\mathrm{\ast,diffuse,max}}$ and $M_{\rm crit}$ as free parameters that we tune to data from the EAGLE simulation for our fiducial model (see Section 3).
We assume that the diffuse stellar component, which describes both the stellar halo as well as the stellar mass of all the satellite galaxies, roughly follows the density distribution of the dark matter. As already mentioned in Section 2.1, this is indeed a good approximation for the radial distribution of satellite galaxies. In addition, both observations (e.g., Montes & Trujillo 2019) and hydrodynamical simulations (e.g., Yoo et al. 2024; Contreras-Santos et al. 2024; Alonso Asensio et al. 2020) have indicated that the stars that make up the stellar halo are a good tracer of the dark matter. We therefore model the density profile of the diffuse stellar component as an NFW profile with a concentration parameter $c_{\mathrm{*,diffuse}}=\eta c_{200}$. For our fiducial model we set $\eta=3$, which yields a good fit to the distribution of diffuse light (ICL) in groups and clusters (Contini & Gu 2020). Note that this approach ignores the fact that satellite galaxies contribute localized density peaks. Rather, the mass of the satellites is modeled as being distributed smoothly following a spherically symmetric profile. As long as the stellar masses of individual satellite galaxies are sufficiently small compared to the host halo, this oversimplification will not have a significant impact on our results. Recall anyways that we assume throughout that satellite galaxies are a kinematic tracer population of the host halo’s gravitational potential.
@@ 2.2.4 Gas
Models of galaxy formation generally predict that more massive halos, with their deeper potential wells, are more effective at retaining baryons (Dekel & Silk 1986), a trend that is also supported by observations (e.g., Dai et al. 2010). We therefore model the ejected baryon fraction, $f_{\rm eject}$, as a decreasing function of halo mass, using a functional form:
$f_{\mathrm{eject}}(M_{\rm DMO})=\frac{f_{\mathrm{eject,max}}}{2}\left[1-{\rm erf}\left(\alpha\,\mathrm{log}\frac{M_{\rm DMO}}{M_{\mathrm{char}}}\right)\right]\,.$ (19)
@@
Hence, the ejected mass fraction transitions from a maximum of $f_{\mathrm{eject,max}}$ at the low mass end, to $f_{\rm eject}=0$ at the high mass end. The steepness of the transition is controlled by $\alpha$, which we treat as a (positive) free parameter. The characteristic mass $M_{\rm char}$ sets the DMO halo mass where $f_{\rm eject}=0.5f_{\rm eject,max}$. The bound gas fraction then automatically follows from equation (18).
@@
We model the bound gas component in halos using a single, smooth, spherically symmetric distribution. In reality, the gas inside dark matter halos consists of multiple components, including the interstellar media (ISM) of the central and satellite galaxies, as well as the circumgalactic medium (CGM) [^1] In the case of massive clusters, the CGM is typically referred to as the intracluster medium (ICM).. However, we focus exclusively on the CGM/ICM component for two main reasons. First, the ISM contributes only a small fraction of the total gas mass (Tumlinson et al. 2017). Second, due to the fibre collision scale (see Section 4.4), our measurements are insensitive to the inner halo regions where the ISM resides. Instead, we primarily probe the diffuse CGM, which extends on scales of order the virial radius. The CGM itself is expected to be multiphase, consisting of cold ($T<10^{4}$ K), cool ($T\sim 10^{4}$$10^{5}$ K), warm ($T\sim 10^{5}$ K) and hot ($T\geq 10^{6}$ K) gas, giving rise to complex features such as clumps, filaments, and streams. Although the hot phase is believed to dominate the CGM mass budget in Milky Way–mass halos, this remains observationally uncertain (Tumlinson et al. 2017). In contrast, at the mass scale of galaxy groups and clusters, X-ray observations provide unambiguous evidence that the hot ($T\sim 10^{7}$$10^{8}$ K) ICM is the dominant component.
@@
We assume the gas is in hydrostatic equilibrium within the gravitational potential of the dark matter halo and follows a polytropic equation of state, $P\propto\rho^{\Gamma}$. Under these assumptions, the gas density profile can be expressed as (e.g. Komatsu & Seljak 2001; Suto et al. 1998; Martizzi et al. 2013):
$\rho_{\rm gas}(x)=\rho_{0}\left[\frac{\ln(1+x)}{x}\right]^{\frac{1}{\Gamma-1}}\,$, (20)
@@
where $x=r/r_{\rm s}$ and $\Gamma$ is the polytropic index. We require that the slope of the gas density profile matches that of the dark matter at a certain transition radius $r_{\rm tr}=x_{\rm tr}r_{\rm s}$, following Komatsu & Seljak (2001). This then fixes the value for $\Gamma$:

### v2

@@ 1 Introduction
The advent of wide-field galaxy surveys has enabled data-driven approaches to empirically infer the galaxy-halo connection and its evolution over time. Popular methods include galaxy clustering (e.g. Berlind and Weinberg 2002; Yang et al. 2003; van den Bosch et al. 2007; Zehavi et al. 2011; Hearin and Watson 2013; Guo et al. 2015b, a, 2016; Zentner et al. 2019), galaxy–galaxy lensing (e.g. Guzik and Seljak 2002; Mandelbaum et al. 2006, 2016; Sonnenfeld and Leauthaud 2018; Leauthaud et al. 2012, 2017), and subhalo abundance matching (SHAM) (Marinoni and Hudson 2002; Conroy et al. 2006; Vale and Ostriker 2004; Reddick et al. 2013). We refer to Wechsler and Tinker (2018) for a comprehensive review of the different approaches to probe and understand the galaxy-halo connection.
@@
A less commonly used technique for probing the galaxy-halo connection is satellite kinematics. This method represents the earliest approach, dating back to the first evidence for dark matter by Zwicky (1933). It involves measuring the line-of-sight velocities of satellite galaxies relative to their central galaxy. Since satellite galaxies are assumed to trace the dark matter potential wells, their dynamics can be used to constrain the gravitational potential and mass of the host halo. Although some concerns have been raised about this method (for an extensive review, see Lange et al. 2019a, and references therein), numerous studies have progressively refined satellite kinematics and demonstrated its utility as a robust tool for constraining the galaxy–halo connection (van den Bosch et al. 2004; Conroy et al. 2007; More et al. 2009b, 2011; Wojtak and Mamon 2013; Lange et al. 2019a, b; Mitra et al. 2024; van den Bosch et al. 2019). In particular, Basilisk, a Bayesian hierarchical inference formalism, introduced by van den Bosch, Lange, and Zentner (2019) and refined by Mitra et al. (2024), has demonstrated exceptional power to probe the galaxy-halo connection using the kinematics of satellite galaxies in large redshift surveys such as the Sloan Digital Sky Survey (SDSS; York et al. 2000). Basilisk  forward models the 2-D projected phase-space distribution of satellite and central galaxy pairs. A key advantage of this forward-modelling approach is its ability to directly compute the likelihood of the full projected phase-space data given the model without the need for stacking the data in luminosity bins. Consequently, it can simultaneously solve for halo mass, the radial profile of the satellite galaxies and orbital anisotropy of the satellite galaxies, while properly accounting for scatter in the galaxy–halo connection. In addition, it properly accounts for biases and selection effects in the data, such as incompleteness, fibre collisions, and interlopers (Lange et al. 2019a; van den Bosch et al. 2019; Mitra et al. 2024). Recently, Mitra et al. (2024) applied Basilisk  to SDSS-DR7 data, yielding inferences on the galaxy–halo connection that are consistent with, yet tighter than, previous constraints derived from other methods like galaxy clustering and galaxy–galaxy lensing, while also overcoming the mass-anisotropy degeneracy and without being sensitive to halo assembly bias.
@@
An important assumption in satellite kinematics studies is the halo’s density profile, which shapes the gravitational potential. Often, such studies, including those conducted with Basilisk, assume that halos are composed solely of dark matter, neglecting the presence of baryons. However, baryons, in the form of gas and stars, can play a critical role in shaping the structure and dynamics of dark matter halos. Processes such as gas cooling, star formation, and feedback from supernovae or active galactic nuclei (AGN) can modify the distribution of both baryons and dark matter. Early models of baryonic effects suggested that gas cooling and dissipation, followed by star formation, can deepen the gravitational potential well (White and Rees 1978; White and Frenk 1991) and lead to adiabatic contraction of dark matter (Blumenthal, Faber, Flores, and Primack 1986). With the advent of cosmological hydrodynamical simulations, detailed studies have further examined the role of baryons in shaping halo structure, clustering, and abundance (e.g. Duffy et al. 2010; Schaller et al. 2015a, b; Sorini et al. 2024; Beltz-Mohrmann and Berlind 2021; Euclid Collaboration et al. 2024). These effects are highly complex, as the influence of baryonic physics depends on factors such as halo mass, environment, and redshift. Moreover, the impact of baryons is extremely sensitive to the specific feedback prescriptions, which differ significantly across hydrodynamical simulations (see review by Somerville and Davé 2015). Despite these challenges, hydrodynamical simulations are often used for correcting for “baryonic effects", that is deviations from dark-matter-only models, when constraining for example the matter power spectrum (e.g. White 2004; van Daalen et al. 2011; Schaller et al. 2025; Schaye et al. 2023; Hernández-Aguayo et al. 2023) and weak lensing observables (Broxterman et al. 2024; Semboloni et al. 2011).
@@
An alternative approach is to quantify baryonic effects analytically, providing a flexible tool to complement hydrodynamical simulations and systematically test their impact on cosmological and large-scale structure inferences. While such an approach has been used to study baryonic effects on the matter power spectrum and weak lensing (Fedeli 2014; Fedeli et al. 2014; Semboloni et al. 2011; Mead et al. 2015; Schneider and Teyssier 2015; Schneider et al. 2019), to our knowledge baryonic corrections have not yet been applied to the inference from satellite kinematics. The main goal of this paper is to rectify this situation and to study how baryons impact the line-of-sight velocity dispersion of satellite galaxies modelled as tracers of the underlying gravitational potential well. We do so by comparing the results of a dark-matter-only (DMO) model with one in which the halo hosts a central galaxy and an extended distribution of diffuse stars and gas. Crucially, we examine how the ejection of baryons from the host halo due to feedback processes impacts the potential well, and consequently, the satellite kinematics. We use these results to develop correction functions that can be used to correct the line-of-sight velocity dispersions of satellites computed using a DMO model for the presence of baryons. We also explore conservative bounds on this baryonic correction model that can be used to marginalize over uncertainties arising from incomplete understanding of physical processes associated with galaxy formation. We specifically tailor these results so they can be used in Basilisk  when applied to SDSS data. In a companion paper (Mitra et al. 2025), it is shown that these baryonic corrections, which are typically only at the level of a few percent, can have an appreciable impact on the inference, affecting both the inferred galaxy halo connection as well as cosmological constraints.
@@
Throughout this paper, we adopt the flat Planck18 CDM cosmology (with BAO constraints, Table 2 in Planck Collaboration VI 2020).
@@ 2.1 Satellite kinematics
$\sigma_{\rm r}^{2}(r)=\frac{G}{r^{2\beta}\,n_{\mathrm{sat}}(r)}\int_{r}^{\infty}r^{\prime 2\beta-2}n_{\mathrm{sat}}(r^{\prime})\,M_{\rm tot}(r)\,\mathrm{d}r^{\prime}\,$, (1)
(e.g. van den Bosch et al. 2004; Łokas and Mamon 2003), where $M_{\rm tot}(r)$ is the total enclosed mass profile, $n_{\mathrm{sat}}(r)$ is the radial number density profile of the satellites, and
@@
Integrating the velocity dispersion in the line-of-sight direction along the line-of-sight, weighted by the number density of the tracer population, yields the line-of-sight velocity dispersion (e.g. Binney and Mamon 1982)
@@
Throughout we follow Mitra et al. (2024) and assume that the radial number density distribution of satellite galaxies follow a generalized Navarro, Frenk, and White (1996) (NFW) profile given by
@@
Finally, we emphasize that the primary observable considered here is the line-of-sight velocity dispersion, i.e. the second moment of the line-of-sight velocity distribution (LOSVD). However, higher-order moments of the LOSVD, such as the skewness and kurtosis, also contain valuable information regarding the kinematics (e.g., van der Marel and Franx 1993). Indeed, Basilisk  uses the kurtosis in its analysis, which helps to break the mass-anisotropy degeneracy (Binney and Mamon 1982). However, as we demonstrate in Appendix B, baryons have a negligible effect on the kurtosis over the radial range of interest, which is why in what follows, we focus solely on the line-of-sight velocity dispersion.
@@
$c_{200}$ Concentration of the halo derived from concentration-halo mass relation. Diemer and Joyce (2019)
@@
$f_{\mathrm{*,diffuse}}$ Fraction of mass in the diffuse stellar component. Best-fit derived from EAGLE: $f_{\mathrm{\ast,diffuse,max}}=0.06$, $\log M_{\mathrm{char,diffuse}}=12.8$
@@
$f_{\mathrm{eject}}$ Mass fraction ejected from halo. Best-fit derived from EAGLE: $f_{\mathrm{eject,max}}=0.66$, $\log M_{\mathrm{char,eject}}=13.2$, $\alpha=1.8$
@@ 2.2 Mass models
The key ingredient for computing $\sigma_{\rm ap}$ is a model for the total enclosed mass, $M_{\rm tot}(r)$, which enters the expression for the radial velocity dispersion given by equation (1). Here we describe the two mass models that we use for our investigation: a DMO model and a baryonic model.
@@
$M_{\rm tot}(r)=$ $M_{\rm DM}(r)+M_{\rm\ast,cen}(r)\,$
$+M_{\rm\ast,diffuse}(r)+M_{\rm gas}(r)\,.$ (9)
We further define $M_{\rm tot}\equiv M_{\rm tot}(r_{200})$ as the total bound mass of the halo in the baryonic model, which is related to the DMO mass according to
@@ 2.2.1 Dark Matter
with $c_{200}=r_{200}/r_{\rm s}$ the halo concentration. Throughout we compute this concentration parameter using the redshift-dependent concentration-mass relation of Diemer and Joyce (2019).
@@ 2.2.2 Central Galaxy
$M_{\mathrm{\ast,cen}}(r)=M_{\mathrm{\ast,cen}}\frac{r^{2}}{(r+a)^{2}}\,.$ (16)
@@ 2.2.3 Diffuse stellar component
Whereas $f_{\mathrm{\ast,cen}}(M_{\rm DMO})$ is tightly constrained via the SHMR, the baryonic mass fraction of the diffuse stellar component, $f_{\mathrm{\ast,diffuse}}$, is poorly constrained. The main reason is that stellar halos typically have extremely low surface brightness, making them difficult to detect. In addition, the outer light profile of the central galaxy typically blends smoothly into that of the stellar halo, hampering a unique decomposition, both in observations (e.g., Huang et al. 2018) and in numerical simulations (e.g., Pillepich et al. 2018). Despite these challenges, it has become clear that $f_{\mathrm{\ast,diffuse}}$ increases with halo mass, albeit with considerable scatter (Merritt et al. 2016; Deason et al. 2019; Ragusa et al. 2023; Montes and Trujillo 2019). We describe this mass dependence using the following simple analytical form:
$f_{\mathrm{\ast,diffuse}}(M_{\rm DMO})=\frac{f_{\mathrm{\ast,diffuse,max}}}{2}\left[1+{\rm erf}\left(\log\frac{M_{\rm DMO}}{M_{\mathrm{char,diffuse}}}\right)\right]\,$, (17)
which transitions from $f_{\mathrm{\ast,diffuse}}=0$ for $M_{\rm DMO}\ll M_{\rm char,diffuse}$ to a maximum of $f_{\mathrm{\ast,diffuse,max}}$ when $M_{\rm DMO}\gg M_{\rm char,diffuse}$. We treat $f_{\mathrm{\ast,diffuse,max}}$ and $M_{\rm char,diffuse}$ as free parameters that we tune to data from the EAGLE simulation for our fiducial model (see Section 3).
We assume that the diffuse stellar component, which describes both the stellar halo as well as the stellar mass of all the satellite galaxies, roughly follows the density distribution of the dark matter. As already mentioned in Section 2.1, this is indeed a good approximation for the radial distribution of satellite galaxies. In addition, both observations (e.g., Montes and Trujillo 2019) and hydrodynamical simulations (e.g., Yoo et al. 2024; Contreras-Santos et al. 2024; Alonso Asensio et al. 2020) have indicated that the stars that make up the stellar halo are a good tracer of the dark matter. We therefore model the density profile of the diffuse stellar component as an NFW profile with a concentration parameter $c_{\mathrm{*,diffuse}}=\eta c_{200}$. For our fiducial model we set $\eta=3$, which yields a good fit to the distribution of diffuse light (ICL) in groups and clusters (Contini and Gu 2020). Note that this approach ignores the fact that satellite galaxies contribute localized density peaks. Rather, the mass of the satellites is modeled as being distributed smoothly following a spherically symmetric profile. As long as the stellar masses of individual satellite galaxies are sufficiently small compared to the host halo, this oversimplification will not have a significant impact on our results. Recall anyways that we assume throughout that satellite galaxies are a kinematic tracer population of the host halo’s gravitational potential.
@@ 2.2.4 Gas
Models of galaxy formation generally predict that more massive halos, with their deeper potential wells, are more effective at retaining baryons (Dekel and Silk 1986), a trend that is also supported by observations (e.g., Dai et al. 2010). We therefore model the ejected baryon fraction, $f_{\rm eject}$, as a decreasing function of halo mass, using a functional form:
$f_{\mathrm{eject}}(M_{\rm DMO})=\frac{f_{\mathrm{eject,max}}}{2}\left[1-{\rm erf}\left(\alpha\,\mathrm{log}\frac{M_{\rm DMO}}{M_{\mathrm{char,eject}}}\right)\right]\,.$ (19)
@@
Hence, the ejected mass fraction transitions from a maximum of $f_{\mathrm{eject,max}}$ at the low mass end, to $f_{\rm eject}=0$ at the high mass end. The steepness of the transition is controlled by $\alpha$, which we treat as a (positive) free parameter. The characteristic mass $M_{\rm char,eject}$ sets the DMO halo mass where $f_{\rm eject}=0.5f_{\rm eject,max}$. The bound gas fraction then automatically follows from equation (18).
@@
We model the bound gas component in halos using a single, smooth, spherically symmetric distribution. In reality, the gas inside dark matter halos consists of multiple components, including the interstellar media (ISM) of the central and satellite galaxies, as well as the circumgalactic medium (CGM) [^1] In the case of massive clusters, the CGM is typically referred to as the intracluster medium (ICM).. However, we focus exclusively on the CGM/ICM component for two main reasons. First, the ISM contributes only a small fraction of the total gas mass (Tumlinson et al. 2017). Second, due to the fibre collision scale (see Section 4.4), our measurements are insensitive to the inner halo regions where the ISM of the centrals resides. Instead, we primarily probe the diffuse CGM, which extends on scales of order the virial radius. The CGM itself is expected to be multiphase, consisting of cold ($T<10^{4}$ K), cool ($T\sim 10^{4}$$10^{5}$ K), warm ($T\sim 10^{5}$ K) and hot ($T\geq 10^{6}$ K) gas, giving rise to complex features such as clumps, filaments, and streams. Although the hot phase is believed to dominate the CGM mass budget in Milky Way–mass halos, this remains observationally uncertain (Tumlinson et al. 2017). In contrast, at the mass scale of galaxy groups and clusters, X-ray observations provide unambiguous evidence that the hot ($T\sim 10^{7}$$10^{8}$ K) ICM is the dominant component.
@@
We assume the gas is in hydrostatic equilibrium within the gravitational potential of the dark matter halo and follows a polytropic equation of state, $P\propto\rho^{\Gamma}$. Under these assumptions, the gas density profile can be expressed as (e.g. Komatsu and Seljak 2001; Suto et al. 1998; Martizzi et al. 2013):
$\rho_{\rm gas}(x)=\rho_{0}\left[\frac{\ln(1+x)}{x}\right]^{1/(\Gamma-1)}\,$, (20)
@@
where $x=r/r_{\rm s}$ and $\Gamma$ is the polytropic index. We require that the slope of the gas density profile matches that of the dark matter at a certain transition radius $r_{\rm tr}=x_{\rm tr}r_{\rm s}$, following Komatsu and Seljak (2001). This then fixes the value for $\Gamma$: